# A Class of Generalized Shift-Splitting Preconditioners for Double Saddle Point Problems

### v3

@@ Abstract
In this paper, we propose a generalized shift-splitting (GSS) preconditioner, along with its two relaxed variants to solve the double saddle point problem (DSPP). The convergence of the associated GSS iterative method is analyzed, and sufficient conditions for its convergence are established. Spectral analyses are performed to derive sharp bounds for the eigenvalues of the preconditioned matrices. Numerical experiments based on examples arising from the PDE-constrained optimization problems demonstrate the effectiveness and robustness of the proposed preconditioners compared with existing state-of-the-art preconditioners.
@@ 1 Introduction
Suppose $n,m$ and $l$ are given positive integers with $n\geq m\geq l.$ Then, we consider the double saddle point problem (DSPP) in the following form [13]:
@@
where $A\in{\mathbb{R}}^{n\times n}$ is symmetric positive definite (SPD) matrix, $B\in{\mathbb{R}}^{m\times n}$ has full row rank, $C\in{\mathbb{R}}^{l\times m}$ and $D\in{\mathbb{R}}^{l\times l}$ is a symmetric positive semidefinite (SPS) matrix. Further, $p\in{\mathbb{R}}^{n},$ $q\in{\mathbb{R}}^{l}$ and $r\in{\mathbb{R}}^{m}$ are known vectors and $x\in{\mathbb{R}}^{n},$ $y\in{\mathbb{R}}^{l}$ and $z\in{\mathbb{R}}^{m}$ are unknown vectors to be determined. Here, $(\cdot)^{T}$ and $\bm{0}$ represent the transpose of a matrix and zero matrix of appropriate size, respectively.
@@
Owing to the broad applicability of the DSPP (1.1), this paper primarily concentrates on its numerical solution. Nevertheless, the large and sparse nature of the saddle point matrix $\mathcal{B}$ makes the iterative methods more favorable than the direct approaches [35]. In particular, Krylov subspace methods, such as the generalized minimal residual (GMRES) method [36], have emerged prominently as one of the most effective and widely used iterative solvers. However, the ill-conditioned nature and poor spectral properties of the coefficient matrix $\mathcal{B}$ lead to a significant slowdown of the Krylov subspace iterative method. This challenge underscores the need for the development of novel, effective, and robust preconditioners to enhance the performance of Krylov subspace iterative methods.
Assuming $\mathbf{A}=\left[\begin{matrix}A&\bm{0}\\ \bm{0}&D\end{matrix}\right],$ $\mathbf{B}=[B\leavevmode\nobreak\ \leavevmode\nobreak\ C^{T}],$ ${\bf f}=[p^{T},q^{T}]^{T},$ ${\bf g}=r,$ ${\bf x}=[x^{T},y^{T}]^{T}$ and ${\bf y}=z,$ the DSPP can be transformed into the two-by-two block saddle point problems [10]. A surge amount of work has been done in recent years to solve two-by-two block saddle point problems, for instance, Uzawa iterative methods [4, 26, 39], successive overrelaxation methods [6, 19], Hermitian and skew-Hermitian methods [3, 5], and so on. However, these methods do not exploit the three-by-three block structure of $\mathcal{B}.$
To leverage the full block structure of $\mathcal{B},$ various preconditioners have been studied in the literature to solve the DSPP (1.1). When $D=\bm{0}$ in (1.1), block diagonal (BD) and tridiagonal type preconditioners [23, 1], shift-splitting (SS) type preconditioners [14], Uzawa methods [24, 22], etc. have been explored. When $D\neq\bm{0},$ BD preconditioners for the DSPP (1.1) have been investigated in [13]. However, the involvement of the Schur complement matrix makes the computation of the BD preconditioner time-consuming and expensive. By splitting the coefficient matrix $\mathcal{B}$ as $\mathcal{B}=\mathcal{B}_{1}+\mathcal{B}_{2},$ where
@@
For the DSPP (1.1) arising from PDE-constrained optimization problem, Rees et al. [34] introduced BD preconditioner $\mathcal{P}_{\text{D}}$ and constrained preconditioner $\mathcal{P}_{\text{C}}.$ Later, a block triangular (BT) preconditioner is developed in [32]. In [40], the authors introduced two types of preconditioners for solving the DSPP: a block-counter-diagonal preconditioner, denoted as $\mathcal{P}_{\mathrm{BCD}}$, and a block-counter-tridiagonal preconditioner, denoted as $\mathcal{P}_{\mathrm{BCT}}$. By using different approximations of Schur complement, preconditioners in BD and BT formats are constructed in [28, 29, 30]. For more research on solving DSPP in the context of PDE-constrained optimization problems, refer to the works [27, 18, 25].
For non-Hermitian linear systems, Bai et al. [7] introduced the SS preconditioner by applying the matrix SS technique to the coefficient matrix. Building on this concept, several SS-type preconditioners have been studied for two-by-two and three-by-three block saddle point problems; see, for example, [14, 15, 16, 37, 2]. However, SS-type preconditioners remain unexplored for DSPP, despite their potential for high efficiency. Motivated by this, in this paper, we introduce a generalized SS (GSS) iterative method and the associated GSS preconditioner for solving the DSPP by implementing the concept of SS to $\mathcal{B}$. Moreover, we derive sufficient criteria for the convergence of the GSS iterative method. The main contributions of the paper are summarized as follows:
@@
  • Finally, numerical experiments are conducted for the DSPP arising from the PDE-constrained optimization problem to demonstrate the effectiveness of the proposed preconditioners.
The outline of the rest of the paper is as follows. Section 2 investigates the solvability conditions of the DSPP (1.1) and properties of the coefficient matrix $\mathcal{B}.$ The GSS iterative process and associated preconditioner are proposed in Section 3. Section 4 investigates the convergence criteria for the proposed GSS iterative method. Two relaxed variants of the proposed GSS preconditioner are presented in Section 5, and the spectral analyses of the corresponding preconditioned matrices are performed. Section 6 deals with the parameter selection strategies of the proposed preconditioners. Experimental results, analyses and discussions of the proposed and existing preconditioners are discussed in Section 7. At the end, Section 8 includes some concluding statements.
Notation. Throughout the paper, we adopt the following notations. Let $\mathbb{C}^{n}$ and ${\mathbb{R}}^{m\times n}$ denote the collection of all $n$-dimensional complex vectors and $m\times n$ real matrices, respectively. The notations $\|A\|_{F},$ $\sigma(A)$ and $tr(A)$ denote the Frobenious norm, spectrum and trace of the matrix $A\in{\mathbb{R}}^{m\times n},$ respectively. For any vector $x\in\mathbb{C}^{n},$ ${x}^{*}$ represents its conjugate transpose. For any matrix $A\in{\mathbb{R}}^{m\times n}$ having real eigenvalues, $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denote the maximum eigenvalue and minimum eigenvalue of $A,$ respectively. For any $z\in\mathbb{C},$ $\mathfrak{R}(z)$ and $\Im(z)$ represent the real and imaginary part of $z,$ respectively.
@@ 2 Solvability conditions for the DSPP and properties of the matrix $\mathcal{B}$
In this section, we provide the solvability conditions on the block matrices $A,$ $B,$ $C$ and $D$ for the system (1.1) and a few important properties of the matrix $\mathcal{B}.$
@@
Let $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ be $SPD$ matrices. If $B$ has full row rank, then the saddle point matrix $\mathcal{B}$ is nonsingular.
@@
$\left\{\begin{array}[]{cc}Ax+B^{T}z=&\bm{0},\\ Dy+Cz=&\bm{0},\\ -Bx-C^{T}y=&\bm{0}.\end{array}\right.$ (2.4)
We first assert that $x=\bm{0}.$ Then from the first equation in (2.4), we obtain $B^{T}z=\bm{0}.$ Since, $B$ has full row rank, this implies $z=\bm{0}.$ Thus the second equation of (2.4) gives $y=\bm{0}$ as $D$ is SPD. This implies ${\bm{u}}=\bm{0}.$ Next, we assert $z=\bm{0}.$ Then from the first and second equation of (2.4), we find that $x=\bm{0}$ and $y=\bm{0},$ respectively, as $A$ and $D$ are SPD. Hence, ${\bm{u}}=\bm{0}.$ Now, we assume that $x\neq\bm{0}$ and $z\neq\bm{0}.$ Then multiplying by $x^{T}$ from the left side of the first equation of (2.4), we obtain
@@
Substituting (2.5) and $Dy=-Cz$ on (2.6), we have $x^{T}Ax+y^{T}Dy=0,$ which is possible only when $x=\bm{0}$ and $y=\bm{0},$ as $A$ and $D$ are SPD matrices. Thus ${\bm{u}}=\bm{0},$ and hence, $\mathcal{B}$ is nonsingular. $\blacksquare$
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be SPD and $D\in{\mathbb{R}}^{l\times l}$ be SPS. If $B$ and $C$ are of full row rank, then $\mathcal{B}$ is nonsingular.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ be SPD matrices. If $B$ and $C$ are of full row rank, then the matrix $\mathcal{B}$ is positive stable, i.e., $\lambda>0$ for all $\lambda\in\sigma(\mathcal{B}).$
@@
$\left\{\begin{array}[]{cc}Au+B^{T}w&=\lambda u,\\ Dv+Cw&=\lambda v,\\ -Bu-C^{T}v&=\lambda w.\end{array}\right.$ (2.10)
@@
$\lambda\|{\bm{u}}\|^{2}=u^{\star}Au+v^{*}Dv+2{i\mkern 1.0mu}\Im(u^{*}B^{T}w+w^{*}Cv).$ (2.11)
Therefore, from (2.11), we obtain
$\mathfrak{R}(\lambda)=\frac{u^{*}Au+v^{*}Dv}{\|{\bm{w}}\|^{2}}.$ (2.12)
From (2.12), we get $\mathfrak{R}(\lambda)\geq 0$ and $\mathfrak{R}(\lambda)=0$ if and only if $u=\bm{0}$ and $v=\bm{0}.$
Next, we show that the vectors $u$ and $v$ can not be zero simultaneously. First, assume that $u=\bm{0}.$ From the first equation in (2.10), it follows that $B^{T}v=\bm{0}$, which leads to $v=\bm{0},$ as $B$ has full row rank. Substituting $u=\bm{0}$ and $v=\bm{0}$ into the third equation of (2.10) yields $u=\bm{0}$. As a result, we obtain $\bm{u}=\mathbf{0}$, which contradicts the assumption that $\bm{u}$ is an eigenvector. Therefore, we conclude that $u\neq\bm{0}$, which in turn implies that $\mathfrak{R}(\lambda)>0$. This completes the proof. $\blacksquare$

### v4

@@ Abstract
In this paper, we propose a generalized shift-splitting (GSS) preconditioner, along with its two relaxed variants to solve the double saddle point problem (DSPP). The convergence of the associated GSS iterative method is analyzed, and sufficient conditions for its convergence are established. Spectral analyses are performed to derive sharp bounds for the eigenvalues of the preconditioned matrices. Numerical experiments based on examples arising from the PDE-constrained optimization problem and the leaky lid driven cavity problem demonstrate the effectiveness and robustness of the proposed preconditioners compared with existing state-of-the-art preconditioners.
@@ 1 Introduction
Suppose $n,m$, and $l$ are given positive integers with $n\geq m\geq l.$ Then, we consider the double saddle point problem (DSPP) in the following form [13]:
@@
where $A\in{\mathbb{R}}^{n\times n}$, $B\in{\mathbb{R}}^{m\times n}$ has full row rank, $C\in{\mathbb{R}}^{l\times m}$ and $D\in{\mathbb{R}}^{l\times l}$ Further, $p\in{\mathbb{R}}^{n},$ $q\in{\mathbb{R}}^{l}$ and $r\in{\mathbb{R}}^{m}$ are known vectors and $x\in{\mathbb{R}}^{n},$ $y\in{\mathbb{R}}^{l}$ and $z\in{\mathbb{R}}^{m}$ are unknown vectors to be determined. Here, $(\cdot)^{T}$ and $\bm{0}$ represent the transpose of a matrix and zero matrix of appropriate size, respectively. In this paper, we consider that the matrices $A$ and $D$ can be both symmetric or nonsymmetric.
@@
Owing to the broad applicability of the DSPP (1.1), this paper primarily concentrates on its numerical solution. Nevertheless, the large and sparse nature of the saddle point matrix $\mathcal{B}$ makes the iterative methods more favorable than the direct approaches [35]. In particular, Krylov subspace methods, such as the generalized minimal residual (GMRES) method [36], have emerged prominently as one of the most effective and widely used iterative solvers. However, the ill-conditioned nature and poor spectral properties of the coefficient matrix $\mathcal{B}$ lead to a significant slowdown of the Krylov subspace method. This challenge underscores the need for the development of novel, effective, and robust preconditioners to enhance the performance of Krylov subspace methods.
Assuming $\mathbf{A}=\left[\begin{matrix}A&\bm{0}\\ \bm{0}&D\end{matrix}\right],$ $\mathbf{B}=[B\leavevmode\nobreak\ \leavevmode\nobreak\ C^{T}],$ ${\bf f}=[p^{T},q^{T}]^{T},$ ${\bf g}=r,$ ${\bf x}=[x^{T},y^{T}]^{T}$ and ${\bf y}=z,$ the DSPP can be transformed into the two-by-two block saddle point problems [10]. A surge amount of work has been done in recent years to solve two-by-two block saddle point problems, for instance, Uzawa iterative methods [4, 27, 39], shift-splitting (SS) methods [15, 16], successive overrelaxation methods [6, 20], Hermitian and skew-Hermitian methods [3, 5], and so on. However, these methods do not exploit the three-by-three block structure of $\mathcal{B}.$
To leverage the full block structure of $\mathcal{B},$ various preconditioners have been studied in the literature to solve the DSPP (1.1). When $D=\bm{0}$ in (1.1), block diagonal (BD) and tridiagonal type preconditioners [1, 24], SS type preconditioners [2, 14], Uzawa methods [25, 23], etc. have been explored. When $D\neq\bm{0},$ BD preconditioners for the DSPP (1.1) have been investigated in [13]. By splitting the coefficient matrix $\mathcal{B}$ as $\mathcal{B}=\mathcal{B}_{1}+\mathcal{B}_{2},$ where
@@
For the DSPP (1.1) arising from PDE-constrained optimization problem, Rees et al. [34] introduced BD preconditioner $\mathcal{P}_{\text{D}}$ and constrained preconditioner $\mathcal{P}_{\text{C}}.$ Later, a block triangular (BT) preconditioner is developed in [33]. In [40], the authors introduced two types of preconditioners for solving the DSPP: a block-counter-diagonal preconditioner, denoted as $\mathcal{P}_{\mathrm{BCD}}$, and a block-counter-tridiagonal preconditioner, denoted as $\mathcal{P}_{\mathrm{BCT}}$. By using different approximations of Schur complement, preconditioners in BD and BT formats are constructed in [29, 30, 31]. For more research on solving DSPP in the context of PDE-constrained optimization problems, refer to [19, 26, 28].
For non-Hermitian linear systems, Bai et al. [7] introduced the SS preconditioner by applying the matrix SS technique to the coefficient matrix. Building on this concept, several SS-type preconditioners have been studied for two-by-two and three-by-three block saddle point problems with nonsymmetric coefficient matrix; see, for example, [14, 15, 16, 37, 2]. However, SS-type preconditioners remain unexplored for DSPP, despite their potential for high efficiency. Motivated by this, in this paper, we introduce a generalized SS (GSS) iterative method and the associated GSS preconditioner for solving the DSPP by implementing the concept of SS technique to $\mathcal{B}$. Moreover, we derive sufficient criteria for the convergence of the GSS iterative method. The main contributions of the paper are summarized as follows:
@@
  • Finally, numerical experiments are conducted for the DSPP arising from the PDE-constrained optimization problem and the leaky lid driven cavity problem to demonstrate the effectiveness of the proposed preconditioners.
The outline of the rest of the paper is as follows. Section 2 investigates the solvability conditions of the DSPP (1.1) and properties of the coefficient matrix $\mathcal{B}.$ The GSS iterative method and associated preconditioner are proposed in Section 3. Section 4 investigates the convergence criteria for the proposed GSS iterative method. Two relaxed variants of the proposed GSS preconditioner are presented in Section 5, and the spectral analyses of the corresponding preconditioned matrices are performed. Section 6 deals with the parameter selection strategies of the proposed preconditioners. Experimental results, analyses and discussions of the proposed and existing preconditioners are discussed in Section 7. At the end, Section 8 includes some concluding statements.
Notation. Throughout the paper, we adopt the following notations. Let $\mathbb{C}^{n}$ and ${\mathbb{R}}^{m\times n}$ denote the collection of all $n$-dimensional complex vectors and $m\times n$ real matrices, respectively. The notations $\|x\|_{2},$ $\|A\|_{F},$ $\sigma(A),$ ${\vartheta(A)}$ and $tr(A)$ denote the $2$-norm, Frobenius norm, spectrum, spectral radius and trace of the matrix $A\in{\mathbb{R}}^{m\times n},$ respectively. For any vector $x\in\mathbb{C}^{n},$ ${x}^{*}$ represents its conjugate transpose. For any matrix $A\in{\mathbb{R}}^{m\times n}$ having real eigenvalues, $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denote the maximum eigenvalue and minimum eigenvalue of $A,$ respectively. For any $x\in\mathbb{C}^{n},$ $\mathfrak{R}(x)$ and $\Im(x)$ represent the real and imaginary part of $x,$ respectively. We denote $A_{H}:=\frac{A+A^{T}}{2}$ and $D_{H}:=\frac{D+D^{T}}{2}$.
@@ 2 Solvability conditions for the DSPP and properties of the matrix $\mathcal{B}$
In this section, we provide the solvability conditions on the block matrices $A,$ $B,$ $C$ and $D$ for the DSPP (1.1) and a few important properties of the matrix $\mathcal{B}.$ The nonsymmetric coefficient matrix $\mathcal{B}$ possesses the following desirable properties, which are crucial in the theoretical analysis of the iterative method and preconditioners designed to solve DSPP (1.1).
@@
Let $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ with $A_{H}$ and $D_{H}$ be positive semidefinite. If $B$ has full row rank, then
  • (i)
    $\mathcal{B}$ is semipositive real: ${\bm{u}}^{T}\mathcal{B}{\bm{u}}\geq 0$ for all ${\bm{u}}\in{\mathbb{R}}^{n+l+m}.$
  • (ii)
    $\mathcal{B}$ is positive semistable: the real part of all eigenvalues of $\mathcal{B}$ is nonnegative.
Proof.
$(i)$ Let $\bm{u}=[x^{T},\,y^{T}\,z^{T}]^{T}\in{\mathbb{R}}^{n+l+m}.$ Then $\bm{u}^{T}\mathcal{B}\bm{u}=x^{T}Ax+y^{T}Dy.$ Therefore, we have $\bm{u}^{T}\mathcal{B}\bm{u}+\bm{u}^{T}\mathcal{B}^{T}\bm{u}=2(x^{T}A_{H}x+y^{T}D_{H}y)$ and hence, $\bm{u}^{T}\mathcal{B}\bm{u}=x^{T}A_{H}x+y^{T}D_{H}y\geq 0,$ as $A_{H}$ and $D_{H}$ are positive semidefinite. Thus, $\mathcal{B}$ is semipositive real.
$(ii)$ Let $\lambda$ be an eigenvalue of $\mathcal{B}$ and ${\bm{u}}=[u^{T},v^{T},w^{T}]^{T}\in{\mathbb{R}}^{n+l+m}$ is the corresponding eigenvector. Then $\bm{u}^{*}\mathcal{B}\bm{u}=\lambda\|\bm{u}\|_{2}$ and $(\bm{u}^{*}\mathcal{B}\bm{u})^{*}=\bar{\lambda}\|\bm{u}\|_{2}.$ Thus
$\mathfrak{R}(\lambda)$ $=\frac{\bm{u}^{*}(\mathcal{B}+\mathcal{B}^{T})\bm{u}}{2\|\bm{u}\|_{2}}$
$=\frac{\mathfrak{R}(\bm{u})^{T}(\mathcal{B}+\mathcal{B}^{T})\mathfrak{R}(\bm{u})^{T}+\mathfrak{I}(\bm{u})^{T}(\mathcal{B}+\mathcal{B}^{T})\mathfrak{I}(\bm{u})^{T}}{2\|\bm{u}\|_{2}}.$
Then, using $(i),$ we have $\mathfrak{R}(\lambda)\geq 0.$blacksquare
Proposition 2.2.
Let $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ be nonsingular matrices with $A_{H}$ and $D_{H}$ positive definite. If $B$ has full row rank, then the saddle point matrix $\mathcal{B}$ is nonsingular.
@@
$\left\{\begin{array}[]{rc}Ax+B^{T}z=&\bm{0},\\ Dy+Cz=&\bm{0},\\ -Bx-C^{T}y=&\bm{0}.\end{array}\right.$ (2.4)
We first assert that $x=\bm{0}.$ Then from the first equation in (2.4), we obtain $B^{T}z=\bm{0}.$ Since, $B$ has full row rank, this implies $z=\bm{0}.$ Thus the second equation of (2.4) gives $y=\bm{0}$ as $D$ is nonsingular. This implies ${\bm{u}}=\bm{0}.$ Next, we assert $z=\bm{0}.$ Then, from the first and second equation of (2.4), we find that $x=\bm{0}$ and $y=\bm{0},$ respectively, as $A$ and $D$ are nonsingular. Hence, ${\bm{u}}=\bm{0}.$ Now, we assume that $x\neq\bm{0}$ and $z\neq\bm{0}.$ Then multiplying by $x^{T}$ from the left side of the first equation of (2.4), we obtain
@@
Substituting (2.5) and $Dy=-Cz$ on (2.6), we have $x^{T}Ax+y^{T}D^{T}y=0,$ and therefore it must be $x^{T}Ax=0$ and $y^{T}D^{T}y=0$, as both the quantities are nonnegative. However, $x^{T}Ax=x^{T}A_{H}x=0$ and $y^{T}D^{T}y=y^{T}D_{H}y=0$, which implies $x=\bm{0}$ and $y=\bm{0},$ since $A_{H}$ and $D_{H}$ are positive definite matrices. Thus, ${\bm{u}}=\bm{0},$ and hence, $\mathcal{B}$ is nonsingular. $\blacksquare$
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be symmetric positive definite (SPD) matrix and $D\in{\mathbb{R}}^{l\times l}$ be symmetric positive semidefinite matrix. If $B$ and $C$ are of full row rank, then $\mathcal{B}$ is nonsingular.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ be such that $A_{H}$ and $D_{H}$ are positive definite matrices. If $B$ and $C$ are of full row rank, then the matrix $\mathcal{B}$ is positive stable, i.e., $\mathfrak{R}(\lambda)>0$ for all $\lambda\in\sigma(\mathcal{B}).$
@@
$\left\{\begin{array}[]{rc}Au+B^{T}w&=\lambda u,\\ Dv+Cw&=\lambda v,\\ -Bu-C^{T}v&=\lambda w.\end{array}\right.$ (2.10)
@@
$\lambda\|{\bm{u}}\|_{2}^{2}=u^{*}Au+v^{*}Dv+2{i\mkern 1.0mu}\Im(u^{*}B^{T}w+v^{*}Cw).$ (2.11)
On the other hand, from $\bm{u}^{*}\mathcal{B}^{T}\bm{u}=\bar{\lambda}\|u\|_{2}^{2},$ we get
$\bar{\lambda}\|{\bm{u}}\|_{2}^{2}=u^{*}A^{T}u+v^{*}D^{T}v-2{i\mkern 1.0mu}\Im(u^{*}B^{T}w+v^{*}Cw).$ (2.12)
By adding (2.11) and (2.12), we obtain
$(\lambda+\bar{\lambda})\|\bm{u}\|_{2}^{2}$ $=u^{*}(A+A^{T})u+v^{*}(D+D^{T})v$
$=\mathfrak{R}(u)^{T}(A+A^{T})\mathfrak{R}(u)+\mathfrak{I}(u)^{T}(A+A^{T})\mathfrak{I}(u)$
$+\mathfrak{R}(v)^{T}(D+D^{T})\mathfrak{R}(v)+\mathfrak{I}(v)^{T}(D+D^{T})\mathfrak{I}(v)$ (2.13)
Therefore, from (2), we obtain
$\mathfrak{R}(\lambda)=\frac{\mathfrak{R}(u)^{T}A_{H}\mathfrak{R}(u)+\mathfrak{I}(u)^{T}A_{H}\mathfrak{I}(u)+\mathfrak{R}(v)^{T}D_{H}\mathfrak{R}(v)+\mathfrak{I}(v)^{T}D_{H}\mathfrak{I}(v)}{\|{\bm{u}}\|^{2}}.$ (2.14)
Since $A_{H}$ and $D_{H}$ are positive definite matrices, from (2.14), we get $\mathfrak{R}(\lambda)\geq 0$ and $\mathfrak{R}(\lambda)=0$ if and only if $u=\bm{0}$ and $v=\bm{0}.$
Next, we show that the vectors $u$ and $v$ can not be zero simultaneously. First, assume that $u=\bm{0}.$ From the first equation in (2.10), it follows that $B^{T}w=\bm{0}$, which leads to $w=\bm{0},$ as $B$ has full row rank. Substituting $u=\bm{0}$ and $w=\bm{0}$ into the third equation of (2.10) yields $v=\bm{0}$. As a result, we obtain $\bm{u}=\mathbf{0}$, which contradicts the assumption that $\bm{u}$ is an eigenvector. Therefore, we conclude that $u\neq\bm{0}$, which in turn implies that $\mathfrak{R}(\lambda)>0$. This completes the proof. $\blacksquare$
To ensure the properties stated in Propositions 2.2-2.4 are satisfied, throughout the remainder of the paper, we assume $A$ and $D$ are nonsingular matrices, where $A_{H}$ and $D_{H}$ are positive definite matrices.